# Geodetic Exploration of the Interior Structure of Enceladus

### v1

@@ Appendix B Derivation of tidal heating partitioning between the shell and the core
$\Im{(\Lambda_{T})}=\Im\bigg{(}\frac{1}{h_{2}}{(k_{2}+1+5\frac{\delta\rho}{\rho}\epsilon)}\bigg{)},$ (B19)
Combining Equations (B8) and (B19), we can rewrite the expression of tidal heating in the shell as:
$\dot{E}_{\text{shell}}=\frac{105}{10}\frac{(\omega R)^{5}}{G}\xi_{2}\,|h_{2}|^{2}\bigg{[}\Im\bigg{(}{\frac{1}{h_{2}}\big{(}k_{2}+1+5\frac{\delta\rho}{\rho}\epsilon\big{)}}\bigg{)}\bigg{]}\,e^{2}.$ (B20)
This formulation provides the tidal heating in the shell by measurements of the real and imaginary parts of the tidal Love numbers ($k_{2}$, $h_{2}$, $\epsilon_{k_{2}}$, $\epsilon_{h_{2}}$), and information on the density structure. The densities of the layers of the body can be constrained using the static gravity field and libration, and are easier to constrain compared to the rheology of the layers as required in Equation (B8). Tidal heating in the core can be computed by subtracting the tidal dissipation in the shell from the total heat. This derivation is general and can be used for all other icy satellites for which the necessary geodetic observations are or will become available.
Appendix C Markov Chain Monte Carlo Internal Structure Inversion
@@
$\mathcal{L}(\textbf{m})\propto\Pi\exp\bigg{(}-\frac{|d_{obs}^{i}-d_{cand}^{i}(\textbf{m})|^{2}}{2\sigma_{i}^{2}}\bigg{)}$ (C1)
@@
where $i$ runs over the observables, m is the vector consisted of the model parameters, and $d_{obs}$ and $d_{cand}$ are observed and computed candidate data in the algorithm, respectively, and $\sigma_{i}$ is data uncertainty.
@@
Appendix D Probability Distribution Function (PDF) of interior model parameters and correlation
@@
Figure 18: Exploration of the key geophysical interior model parameters corresponding to with Figure 4. Multiple peaks observed in the PDFs of the viscosities of ice shell and the core correspond to separate tidal dissipation regimes that can match the inferred global heat loss. These peaks, and the correlation observed between the viscosities of the shell and rock correspond to the regions indicated in Figure 1. Shear moduli are presented in GPa and other parameters are presented in their SI units.
@@
The sensitivity of the adopted set of measurements used in Figure 18 to parameters such as the density and rigidity of the layers is small. Multiple peaks are apparent in the PDFs of the viscosities of the core and the shell, which are associated with the multiple regions highlighted in Figure 1 that are consistent with the estimated heat loss. These distinct peaks are indicative of tidal dissipation dominated by viscous deformation versus anelastic deformation. In other words, these peaks correspond to viscosities that are lower than or higher than the viscosity associated with the maximum heat production rate.
@@
When considering the constraints on the geodetic observables of Park et al. (2024) and separating the interior parameters for the two tidal dissipation scenarios, we obtain constraints on the interior parameters and correlations between them, as shown in Figures 21 and 22.
@@
Figure 19: Exploration of the correlation between the model parameters associated with Figure 5. Shear moduli are presented in GPa and other parameters are presented in their SI units.
Figure 20: Exploration of the correlation between the model parameters associated with Figure 6. Shear moduli are presented in GPa and other parameters are presented in their SI units.
@@
Figure 21: model parameters for the case where at least 90% of the tidal heating is focused in ice shell. The tidal heating in the shell is calculated using the formulation by (Beuthe, 2019) (Equation (B8)) (Supplementary material figure).
@@
Figure 22: Model parameters for the case where at least 90% of the tidal heating is focused in the rocky core. The tidal heating in the shell is calculated using the formulation in (Beuthe, 2019) (Equation (B8)) and the heating in the core is calculated by subtracting the heat in the shell from the total tidal heating computed based on Equation (5) (Supplementary material figure).

### v3

@@ Appendix C Derivation of tidal heating partitioning between the shell and the core
$\Im{(\Lambda_{T})}=\Im\bigg{(}\frac{1}{h_{2}}{(k_{2}+1+5\frac{\delta\rho}{\rho}\epsilon)}\bigg{)},$ (C19)
Combining Equations (C8) and (C19), we can rewrite the expression of tidal heating in the shell as:
$\dot{E}_{\text{shell}}=\frac{105}{10}\frac{(\omega R)^{5}}{G}\xi_{2}\,|h_{2}|^{2}\bigg{[}\Im\bigg{(}{\frac{1}{h_{2}}\big{(}k_{2}+1+5\frac{\delta\rho}{\rho}\epsilon\big{)}}\bigg{)}\bigg{]}\,e^{2}.$ (C20)
This formulation provides the tidal heating in the shell by measurements of the real and imaginary parts of the tidal Love numbers ($k_{2}$, $h_{2}$, $\epsilon_{k_{2}}$, $\epsilon_{h_{2}}$), and information on the density structure. The densities of the layers of the body can be constrained using the static gravity field and libration, and are easier to constrain compared to the rheology of the layers as required in Equation (C8). Tidal heating in the core can be computed by subtracting the tidal dissipation in the shell from the total heat. This derivation is general and can be used for all other icy satellites for which the necessary geodetic observations are or will become available.
Appendix D Markov Chain Monte Carlo Internal Structure Inversion
@@
$\mathcal{L}(\textbf{m})\propto\Pi\exp\bigg{(}-\frac{|d_{obs}^{i}-d_{cand}^{i}(\textbf{m})|^{2}}{2\sigma_{i}^{2}}\bigg{)}$ (D1)
@@
where $i$ runs over the observables, m is the vector composed of the model parameters, and $d_{obs}$ and $d_{cand}$ are observed and computed candidate data in the algorithm, respectively, and $\sigma_{i}$ is data uncertainty.
@@
Appendix E Probability Density Function (PDF) of interior model parameters and correlation
@@
Figure 18: Exploration of the key geophysical interior model parameters corresponding to with Figure 4. Multiple peaks observed in the PDFs of the viscosities of ice shell and the core correspond to separate tidal dissipation regimes that can match the inferred global heat loss. These peaks, and the correlation observed between the viscosities of the shell and rock correspond to the regions indicated in Figure 1. Shear moduli are presented in GPa, $\rm D_{shell}$ and $\rm R_{core}$ are in km, and other parameters are presented in their SI units. Ice shell thickness, and shear moduli are plotted in logarithmic axes.
@@
Figure 18 shows corner plots of the interior parameters associated with the case considered in Figure 1. Here, the sensitivity of the adopted set of measurements used in Figure 18 to parameters such as the density and rigidity of the layers is small. In addition, in this case we have assumed a very broad range of viscosities of the ice shell and core. As a result, multiple peaks are apparent in the PDFs of the viscosities of the core and the shell, which are associated with the multiple regions highlighted in Figure 1, consistent with the estimated heat loss. These distinct peaks are indicative of tidal dissipation dominated by viscous deformation versus anelastic deformation. In other words, these peaks correspond to viscosities that are lower than or higher than the viscosity associated with the maximum heat production rate.
Assuming the same set of measurements, but limiting the viscosity ranges of the shell and core to $10^{11}<\eta_{shell}<10^{15}$  Pa.s and $10^{11}<\eta_{core}<10^{20}$ Pa.s, and separating the two sets of models based on the two dissipation scenarios, we obtain the interior parameters shown in Figures 17a and 17b.
We compare the amplitude and phase lags of $h_{2}$ and $l_{2}$ in Figure 19 for the two tidal dissipation scenarios. The 2-$\sigma$ standard deviations are indicated in this figure.
Figure 19: Vertical and horizontal displacement Love numbers ($|h_{2}|$, $|l_{2}|$) and phase lags ($\epsilon_{h_{2}}$, $\epsilon_{l_{2}}$) are presented for the core-dominated and shell-dominated tidal dissipation. The heat flux is assumed to be $\rm 15~GW<\dot{E}<40~GW$, covering estimates in Hemingway et al. (2018) and Park et al. (2024), and the direct measurements (Howett et al., 2011). Libration is conservatively assumed to be between $\rm 0.082^{\circ}<\gamma<0.169^{\circ}$, covering the estimates in Thomas et al. (2016), Nadezhdina et al. (2016), and Park et al. (2024). The top axis in the plots for $h_{2}$ and $l_{2}$ indicate the associated maximum radial and maximum easterly displacements corresponding to the values of the Love number in the bottom axis, respectively.
@@
When considering the constraints on the geodetic observables of Park et al. (2024) and separating the interior parameters for the two tidal dissipation scenarios, we obtain constraints on the interior parameters and the correlations between them, as shown in Figures 18b and 18c.
Figure 17a: Exploration of the correlation between the model parameters associated with Figure 5 for the case tidal dissipation is dominated in the shell. Black and yellow colors indicate lowest and highest likelihoods. Units and plotting conventions are similar to Figure 18.
@@
Figure 17b: Exploration of the correlation between the model parameters associated with Figure 5 for the case tidal dissipation is dominated in the shell. Units and plotting conventions are similar to Figure 18.
Figure 18a: Exploration of the correlation between the model parameters associated with Figure 7. The ranges of the plausible interior parameters are shown in Table 4. Blue and red colors indicate the lowest and highest likelihoods. Units and plotting conventions are similar to Figure 18.
@@
Figure 18b: Model parameters for the case where at least 90% of the tidal heating is focused in ice shell associated with Figures 7 and 18a. The tidal heating in the shell is calculated using the formulation by (Beuthe, 2019) (Equation (C8)). Black and yellow colors indicate lowest and highest likelihoods. Units and plotting conventions are similar to Figure 18.
@@
Figure 18c: Model parameters for the case where at least 90% of the tidal heating is focused in the rocky core associated with Figures 7 and 18a. The tidal heating in the core is calculated by subtracting the heat in the shell from the total tidal heating computed based on Equation (5). Blue and red colors indicate the lowest and highest likelihoods. Units and plotting conventions are similar to Figure 18.
Appendix F Effect of shell heterogeneous structure on the Love numbers
Here, we assess the effect of lateral variations in ice shell thickness on the tidal response and its effect on the required measurements of $k_{2m}$ and $h_{2m}$. We use the shape model shown in Figure 11.We build on the analysis of Berne et al. (2023a) and conduct simulations of tidal deformation for four models: 1. without lateral variations in ice shell thickness; 2. with lateral variations in ice shell thickness; 3. with a dissipative core and without lateral variations in ice shell thickness; and 4. with a dissipative core and lateral variations in ice shell thickness. For each model class, we vary the mean ice shell thickness across 15–50 km consistent with the findings in Section 6.1. We assume a core of viscosity $10^{20}$ Pa.s for the case of a non-dissipative core, and $10^{12}$ Pa.s for the case of a dissipative core. All other parameters are chosen to be the same as those used in Figures 1 and 2. To compute $k_{2m}$ and $h_{2m}$, we use a numerical finite element model (Berne et al., 2023a).
Figure 19: Top: Gravitational Love numbers, $k_{20}$ (left column) and $k_{22}$ (right column), against mean ice shell thickness for models that assume an ice shell of constant thickness (dotted line), an ice shell with lateral variations in thickness (solid line), or a dissipative core (dotted and solid red lines). Bottom: Similar to top row except showing displacement Love numbers $h_{20}$ and $h_{22}$ plotted against mean ice shell thickness. x-axes are plotted on a log-scale.
If the core is tidally active and is the primary location of the tidal heating, then its effect of $k_{2m}$ is very significant, regardless of the shell thickness. In that case $k_{2m}$ increases by a factor of (2–4). If the shell thickness is small, the effect of structural heterogeneity on both $k_{2m}$ and $h_{2m}$ is important. For a thin shell, $h_{20}$ can vary by a maximum of 10% while $h_{22}$ can vary by approximately 50%, as a result of the shell thickness variation. For thick shells, the effects of the tidal activity of the core and the structural heterogeneity of the shell on $h_{2m}$ are small because the tidal displacement is dominated by the shell rigidity. For a thin shell, the effects of the structural heterogeneity of the shell and the tidal activity of the core on $k_{2m}$ become superimposed. However, the effect of core activity is more significant. If the tidal activity is focused in the core, then its effect on the shell displacement ($h_{2m}$) through the additional gravitational perturbation can be more significant than the shell structural heterogeneity and overshadow it.